OpenStudy (anonymous):

How do I graph this function? y=1/2|x|

OpenStudy (phi):

break it into 2 equations y = 1/2 x , x>=0 y = -1/2 x, x<=0 now pick some x values for example -2,-1, 0 , 1, 2 and plot the points (x,y)

OpenStudy (anonymous):

You can make a table for these types of problems: so